A labourer is engaged for 20 days on the condition that he will receive Rs 60 for each day he works and will be fined Rs 5 for each day he is absent. If receives Rs 745 in all, for how many days did he remain absent?
Answer by Earlsdon(6294)

Let W = the number of days the labourer worked and A = the number of days the labourer was absent.
Given:
1) W+A = 20 "A labourer was engaged for 20 days..."
2) 60W-5A = 745 "...will receive (+) Rs 60 for each day he works and will be fined (-) Rs 5 for each day he is absent."
Rewrite equation 1) as:
1a) A = 20-W and substitute into equation 2) for A
2a) 60W-5(20-W) = 745 Simplify and solve for W.
2b) 60W-100+5W = 745
2c) 65W = 845
2d) W = 13 and A = 20-13 = 7.
The labourer was absent for 7 days.